how do i write equation y=...form
5(x+y)=20+3x

Answer by checkley75(3666)

5(x+y)=20+3x
5x+5y=20+3x
5y=3x-5x=20
5y=-2x+20
y=-2x/5+20/5
y=-2x/5+4